OpenStudy (anonymous):

4 6x-4 2+ -- = ----- x+2 x^2-4

OpenStudy (kc_kennylau):

Factorize the numerator and the denominator of the second fraction.

OpenStudy (anonymous):

@kc_kennylau thank you so much for all the help like honestly

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

4 2(3x-2) 2+ -- = ------- x+2 (x-2) (x+2)

OpenStudy (kc_kennylau):

Make 2 a fraction with denominator being x+2.\(\left(a=\dfrac {ax}x\right)\)

OpenStudy (anonymous):

2(x+2) (x-2) ---------- (x+2)(x-2)

OpenStudy (kc_kennylau):

you mean: 2(x+2) (x-2) 4 2(3x-2) ----------+---=---------- (x+2)(x-2) x+2 (x-2)(x+2)

OpenStudy (anonymous):

does it come out to be 2( x+2) (x-2)+4(x-2)=2x3(x-2)(x+2) ------------ (x+2) (x-2)

OpenStudy (kc_kennylau):

no

OpenStudy (kc_kennylau):

It comes out to be 2+4(x-2)=2(3x-2)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

how did you get that

OpenStudy (kc_kennylau):

2(x+2) (x-2) 4 2(3x-2) ----------+---=---------- (x+2)(x-2) x+2 (x-2)(x+2) 2(x+2) (x-2) 4(x-2) 2(3x-2) ----------+---------=---------- (x+2)(x-2) (x+2)(x-2) (x-2)(x+2) 2(x+2)(x-2)+4(x-2)=2(3x-2) Sorry for my mistake.

OpenStudy (anonymous):

why do you loose the denominator in the last step?

OpenStudy (kc_kennylau):

I multiplied the whole equation by (x+2)(x-2).

OpenStudy (anonymous):

which side of the equation

OpenStudy (kc_kennylau):

both.

OpenStudy (kc_kennylau):

so that the equality is still valid.

OpenStudy (anonymous):

oh then i did this completely wrong

OpenStudy (anonymous):

ok so would it be one of those fractional equations that you multiply the denominator toboth sides

OpenStudy (anonymous):

ok so i got 2(x+2)(x-2)+4(x-2)=2(3x-2)

OpenStudy (kc_kennylau):

yep, that's the same with mine

OpenStudy (anonymous):

okay thanks so much